OpenStudy (anonymous):

The equation sin (pi/3 -x) is equal to ?

OpenStudy (anonymous):

sin(pi/3 - x) = sin(pi/3)cosx - sinxcos(pi/3) = (sqrt3/2)cosx - sinx(1/2)

OpenStudy (anonymous):

= (sqrt3)/2*cosx - (1/2)*sinx

OpenStudy (anonymous):

ah sorry i forgot to say: a. 1/2(√3cos x+ sin x) b.1/2(√3 cos x-sin x) c. 1/2(cos x-√3 sin x) d. 1/2(cos x+ √3 sin x)

OpenStudy (nali):

still need help?